# Team Samsung-RAL: Technical Report for 2024 RoboDrive Challenge-Robust Map Segmentation Track

### v1

@@ 2.1 Semantic Map Construction
Semantic map construction is a prominent and extensively researched area within the field of autonomous driving. HDMapNet [8] first introduces the problem of local semantic map learning, which aims to construct semantic maps online from the observations of LiDAR sensors and cameras. It also proposes a learning method to build BEV features from sensory input and predicts vectorized map elements. BEVSegFormer [12] proposes the multi-camera deformable attention to transform image-view features to BEV representations for semantic map construction. Very recently, BEVerse [18] incorporates the semantic map construction as part of the multitask framework and uses vanilla convolutional layers for segmentation prediction. MBFusion [3] proposes a multi-modal BEV feature fusion approach for the HD map construction task. In addition, high-definition maps (HD maps) have received widespread attention, which can provide fine-grained information about road scenes and play a vital role in autonomous vehicles.
@@ 2.2 Robustness Under Natural Corruptions
Recently, assessing autonomous driving perception model robustness under natural corruptions has burgeoned as a pivotal research domain [2, 23, 6, 16, 19]. The natural corruption robustness often refers to the capability of a conventionally trained model maintaining satisfactory performance under natural distribution shifts. ModelNet-C [13] design a taxonomy of common 3D corruptions and identify the atomic corruptions to understand classifiers’ robustness. Recently, the corruption robustness of BEV perception tasks has been widely studied. Robodepth [6, 13] establishes a robustness benchmark for monocular depth estimation under corruptions. RoboBEV [16] introduces a comprehensive benchmark for evaluating the robustness under natural corruptions of four BEV perception algorithms, such as 3D object detection [9, 10], semantic segmentation [18, 21], depth estimation [14], and semantic occupancy prediction [15, 4]. [2] systematically designs 27 types of common corruptions in 3D object detection for both LiDAR and camera sensors. Robo3D [5] presents a comprehensive benchmark for probing the robustness of 3D detectors and segmentors under out-of-distribution scenarios against natural corruptions that occur in real-world environments. In this technical report, we study the robustness of semantic map segmentation models under sensor corruption.

### v2

@@ 2.1 Semantic Map Construction
Semantic map construction is a prominent and extensively researched area within the field of autonomous driving. HDMapNet [11] first introduces the problem of local semantic map learning, which aims to construct semantic maps online from the observations of LiDAR sensors and cameras. It also proposes a learning method to build BEV features from sensory input and predicts vectorized map elements. BEVSegFormer [15] proposes the multi-camera deformable attention to transform image-view features to BEV representations for semantic map construction. Very recently, BEVerse [21] incorporates the semantic map construction as part of the multitask framework and uses vanilla convolutional layers for segmentation prediction. MBFusion [5] proposes a multi-modal BEV feature fusion approach for the HD map construction task. In addition, HD map construction [4, 3] have received widespread attention, which can provide fine-grained information about road scenes and play a vital role in autonomous vehicles.
@@ 2.2 Robustness Under Natural Corruptions
Recently, assessing autonomous driving perception model robustness under natural corruptions has burgeoned as a pivotal research domain [2, 26, 8, 19, 22, 10]. The natural corruption robustness often refers to the capability of a conventionally trained model maintaining satisfactory performance under natural distribution shifts. ModelNet-C [16] design a taxonomy of common 3D corruptions and identify the atomic corruptions to understand classifiers’ robustness. Recently, the corruption robustness of BEV perception tasks has been widely studied. Robodepth [8, 16] establishes a robustness benchmark for monocular depth estimation under corruptions. RoboBEV [19] introduces a comprehensive benchmark for evaluating the robustness under natural corruptions of four BEV perception algorithms, such as 3D object detection [12, 13], semantic segmentation [21, 24], depth estimation [17], and semantic occupancy prediction [18, 6]. [2] systematically designs 27 types of common corruptions in 3D object detection for both LiDAR and camera sensors. Robo3D [7] presents a comprehensive benchmark for probing the robustness of 3D detectors and segmentors under out-of-distribution scenarios against natural corruptions that occur in real-world environments. In this technical report, we study the robustness of semantic map segmentation models under sensor corruption.